Q77 / 160 · AP EAMCET 2019 (22 Apr Shift 1) · Mathematics
The probability of occurrence of an event is \(\frac{2}{5}\) and the probability of non-occurrence of another event is \(\frac{3}{10}\). If these events are independent, then the probability that only one of the two events occur is
- A\(\frac{27}{25}\)
- Correct answerB\(\frac{27}{50}\)
- C\(\frac{7}{25}\)
- D\(\frac{14}{25}\)
Solution
Given,
\(P(A)=\frac{2}{5}\)
\(\begin{aligned}
& \therefore \quad P(A)^{\prime}=1-P(A)=1-\frac{2}{5}=\frac{3}{5} \\
& \text { and } \\
& P(B)^{\prime}=\frac{3}{10} \\
& \therefore \quad P(B)=1-P(B)^{\prime}=1-\frac{3}{10}=\frac{7}{10} \\
& \text { Required probability }=P(A) P(B)^{\prime}+P(A)^{\prime} P(B) \\
& =\frac{2}{5} \times \frac{3}{10}+\frac{3}{5} \times \frac{7}{10} \\
& =\frac{6}{50}+\frac{21}{50}=\frac{27}{50}
\end{aligned}\)
Hence, the correct answer is option: 'B'.